Activity:



Find CT, V1, V2, V3 & UT, when C1 = 1F, C2 = 2F, C3 = 3F & QT = 4C.




1
Expert's answer
2022-04-06T13:57:20-0400

"\frac{1}{C_T}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}"

"\frac{1}{C_T}=\frac{1}{1}+\frac{1}{2}+\frac{1}{3}"

"C_T=\frac{6}{11}F"

"Q_T=Q_1=Q_2=Q_3=4C"

"V_1=\frac{Q}{C}=\frac{4}{1}=4V"

"V_2=\frac{Q}{C}=\frac{4}{2}=2V"

"V_3=\frac{Q}{C}=\frac{4}{3}=1.33V"


"U_T=\frac{1}{2}\frac{Q^2}{2C_T}=\frac{1}{2}\times\frac{4^2\times11}{6}=29.33J"